# LetsGo: Large-Scale Garage Modeling and Rendering via LiDAR-Assisted Gaussian Primitives

### v1

@@ Abstract
Large garages are ubiquitous yet intricate scenes in our daily lives, posing challenges characterized by monotonous colors, repetitive patterns, reflective surfaces, and transparent vehicle glass. Conventional Structure from Motion (SfM) methods for camera pose estimation and 3D reconstruction fail in these environments due to poor correspondence construction. To address these challenges, this paper introduces LetsGo, a LiDAR-assisted Gaussian splatting approach for large-scale garage modeling and rendering. We develop a handheld scanner, Polar, equipped with IMU, LiDAR, and a fisheye camera, to facilitate accurate LiDAR and image data scanning. With this Polar device, we present a GarageWorld dataset consisting of five expansive garage scenes with diverse geometric structures and will release the dataset to the community for further research. We demonstrate that the collected LiDAR point cloud by the Polar device enhances a suite of 3D Gaussian splatting algorithms for garage scene modeling and rendering. We also propose a novel depth regularizer for 3D Gaussian splatting algorithm training, effectively eliminating floating artifacts in rendered images, and a lightweight Level of Detail (LOD) Gaussian renderer for real-time viewing on web-based devices. Additionally, we explore a hybrid representation that combines the advantages of traditional mesh in depicting simple geometry and colors (e.g., walls and the ground) with modern 3D Gaussian representations capturing complex details and high-frequency textures. This strategy achieves an optimal balance between memory performance and rendering quality. Experimental results on our dataset, along with ScanNet++ and KITTI-360, demonstrate the superiority of our method in rendering quality and resource efficiency.
Figure 1: We present LetsGo - an explicit and efficient end-to-end framework for high-fidelity rendering of large-scale garages. We design a handheld Polar scanner to capture RGBD data of expansive parking environments and have scanned a garage dataset, named GarageWorld, comprising five garages with different structures. Our LiDAR-assisted Gaussian primitives approach along with GarageWorld dataset enables various applications, such as autonomous vehicle localization, navigation and parking, as well as VFX production.
**footnotetext: These authors contributed equally to this work.footnotetext: These authors are corresponding authors.
@@ 1 Introduction
For many of us, our daily life begins with a safe departure from a garage and ends with a safe arrival. The garage serves as the origin of our journey to innovation. For visual artists, garages represent a frontier in visual simulation that merges the aesthetic with the technical, offering a canvas where the intricacies of light, shadow, texture, and space coalesce. The complex interplay of artificial and natural lighting within the confines of a garage, with its reflective surfaces, varying materials, and intricate geometries, provides a rigorous testbed for modeling and rendering, pushing the boundaries of what is achievable in virtual environments.
A number of cutting-edge fields can also readily benefit from real-world garage modeling and rendering. For autonomous driving, high-fidelity 3D models of garages are crucial for simulating and testing the complex navigation and parking algorithms that self-driving vehicles must master. By rendering real-world garages with photorealistic details, developers can create varied scenarios, including different obstacles, ensuring that autonomous systems can operate safely and efficiently in real environments. In game design, 3D garages serve as interactive backdrops that can be rich in narrative and aesthetic detail. Digital replications of real-world garages provide a sense of immersion, crucial for players to explore and interact with. For the movie industry, the ability to render photorealistic garages allows for visually stunning special effects that can be seamlessly integrated into live-action footage. This capability is invaluable for creating believable scenes in a controlled environment without the logistical challenges and costs associated with on-location shooting.
@@
Large-scale garages in the real world, however, are inherently difficult to model. They are generally poorly lit and are mainly composed of texture-less walls, sloped surfaces with distorted lines, cluttered and often reflective environments, and complex occlusions between walls, cars, and pillars (Fig. 11). Garages with internal circular or spiral paths further necessitate sophisticated handling of the continuous curvature and potential occlusions caused by the spiral design, which can lead to incomplete data and ambiguities in spatial relationships. Furthermore, the extensive spatial domain of these garages underscores the necessity for efficient 3D representations and lightweight rendering techniques to facilitate real-time interaction and visualization, particularly for scenarios demanding rapid situational assessment, e.g., in navigation and path adjustment.
@@
Traditional computer vision techniques have long struggled to reconstruct accurate 3D models of garages. Passive perception schemes primarily employ color cameras to capture images, utilizing Structure from Motion (SfM) [56, 62, 60, 61] and Multi-view Stereo (MVS) [12, 22, 77, 75, 74] for reconstructing 3D geometry. However, due to the prevalence of texture-less regions and repetitive structural designs within garages, SfM and MVS methods often fail to extract sufficient feature points and establish accurate feature correspondences necessary for estimating camera poses. Active sensing technologies based on LiDAR can calculate camera poses and scene geometry using SLAM [29, 15, 32, 6] algorithms, but the reflective materials and transparent car windows common in garages lead to geometric inaccuracies. More importantly, the LiDAR data tend to be sparse, where the results contain many holes that corrupt high-frequency textures essential for rendering the color appearance of the scene.
@@
Recent advances in neural representation, particularly the Neural Radiance Fields (NeRF) [40] approach, have shown promise in producing high-quality renderings given camera parameters, but they come with high computational costs and lengthy training times. A succession of NeRF-based [44, 9, 79, 26, 50, 82] enhancements have emerged to optimize training duration and visual rendering quality. However, integrating such implicit representations into conventional graphics rendering pipelines and tools for rapid 3D content applications remains challenging. The emerging 3D Gaussian Splatting [25] (3DGS) method has revisited explicit representations, utilizing 3D Gaussians to articulate the geometry and appearance of scenes. This explicit modality not only achieves high-quality scene modeling and rendering but also integrates seamlessly into existing 3D content production workflows. Nonetheless, 3DGS requires a relatively accurate point cloud from SfM for Gaussian initialization, while SfM often fails in the large-garage environment with large texture-less regions and high similarity between different parts of the garages. Furthermore, 3DGS requires a substantial number of 3D Gaussians to depict even the simplest geometries like walls and floors, leading to considerable memory and storage resource consumption that significantly hinders its application in large-scale scenes.
@@
This paper introduces LetsGo - an explicit and efficient end-to-end modeling scheme for high-fidelity rendering of large-scale garages. Our critical insight is to aid the 3D Gaussian splatting variants with calibrated LiDAR points. We design a handheld Polar scanner for expansive garage data collection, which combines IMU, LiDAR and a fisheye camera for robust relative pose estimation. We have scanned five large-scale garages, named GarageWorld, with different geometric structures with this Polar scanner. To our knowledge, this dataset is the first aimed at large-scale garages and will be released to the community. We demonstrate that the LiDAR points collected by this Polar device successfully assist a suite of Gaussian splatting algorithms for garage scene representation. To improve the quality of the 3D Gaussian rendering, we also introduce a depth regularizer that uses depth priors as supervisory signals, significantly reducing floating artifacts and enabling high rendering quality.
@@
Additionally, we develop a lightweight web renderer that supports Level of Detail (LOD) rendering for 3D Gaussians based on the camera’s position, orientation, and frustum, enabling real-time high-quality rendering of large-scale scenes on various consumer-level devices. Our results, gathered from our dataset as well as from ScanNet++ [11] and the KITTI-360 [33] datasets, indicate that our approach not only exceeds other methods in rendering quality but also maintains a smaller memory footprint. Our GarageWorld dataset and the LiDAR-assisted Gaussian primitives for large-scale garage modeling and rendering enables various applications (Fig. 1), including autonomous driving, localization, navigation, and visual effects, etc.
Our primary contributions are as follows:
  • We design a handheld Polar scanner with calibrated IMU, LiDAR, and a fisheye camera, allowing scanning and reconstruction of large-scale garage scenes with large texture-less regions and repetitive patterns where conventional SfM fails.
  • We present the first garage dataset, GarageWorld, consisting of five large-scale parking scenes with diverse structures, which will be released to the community to tackle modeling and rendering challenges in expansive environments.
  • We demonstrate our LiDAR points assist a suite of Gaussian splatting algorithms, making them feasible for expansive and challenging garage scene modeling. We also introduce a depth prior for Gaussian primitive representation training, and a lightweight renderer to enable real-time rendering on consumer-level devices.
@@ 2 Related Work
Conventional Explicit Visual Reconstruction. Conventional algorithms for reconstructing large-scale scenes include Structure from Motion (SfM) [63, 10, 71, 65, 42], Simultaneous Localization and Mapping (SLAM) [29, 4, 5, 7] and Multi-View Stereo (MVS) [58, 18]. They are dedicated to discerning the three-dimensional structure of a scene through the sequential or multi-view analysis of two-dimensional image frames. All these methods leverage feature tracking and multi-view consistency to recover the 3D scene structures. SfM- and SLAM-based methods [63, 10, 56, 1, 13, 21, 67, 55, 31] estimate poses of input images and recover the scene structure jointly. However, their main purpose is pose estimation, and the recovered scene point clouds are always sparse, making it difficult for high-quality free-view synthesis. While MVS-based methods [58, 18, 30, 12, 22, 77, 75, 74] require posed images as input. It computes a dense depth map for each input image, and thus, the constructed point clouds for 3D scene representation are dense. Despite successes, the constructed scenes often lack accuracy and robustness in texture-less and complex scenes.
Implicit Neural Scene Representations. In recent years, neural scene representation [44, 9, 79, 26, 50] has emerged as a promising avenue for addressing the complexities of novel view synthesis. Unlike traditional explicit representation methods, neural approaches leverage the power of deep learning to learn implicit representations directly from data. These methods have shown significant advancements in capturing complex 3D scenes and generating novel views with remarkable realism.
Neural Radiance Fields (NeRF), initially introduced by Mildenhall et al. [40], has revolutionized the field of 3D reconstruction and provided an innovative framework for capturing complex scene details. Since then, a multitude of research has flourished. Considering the original NeRF can only handle bounded and forward-facing scenes, NeRF++ [82] firstly extends NeRF to unbounded scenes by introducing an inverted sphere parameterization. MipNeRF [2] reduces objectionable aliasing artifacts when training and testing images are at different scene resolutions by introducing a new conical frustum rendering scheme instead of rendering along a camera ray. Based on MipNeRF, MipNeRF360 [3] further leverages a non-linear scene representation, online distillation, and a novel distortion-based regularizer to improve the rendering quality in unbounded scenes. NeRF has also been extended to address deformable scenes [46, 38, 37] and scenes with dynamic objects [81, 49, 47].
The original NeRF requires long training and evaluation time, prohibiting its practical use. To achieve real-time 3D modeling and rendering, kiloNeRF [52] proposes to leverage many tiny MLPs to replace the original big MLP, as tiny MLPs require a significantly shorter time for evaluation than the original large MLP. Instant-NGP [43] introduces a new hash encoding approach, which guarantees rendering quality with a smaller neural network and thus reduces the rendering speed. To handle large-scale scenes, Block-NeRF [66] decomposes the scene into individually trained NeRFs, which decouples the rendering time from scene size and enables per-block updates of the environment. Turki et al. [68] introduce Mega-NeRF, which handles varying lighting conditions of images spanning different buildings or city blocks and explores temporal consistency. Considering that previous NeRFs are mainly designed for forward-facing or the 360 object-centric trajectory, F2-NeRF [70] presents a novel space-warping method to handle arbitrary camera trajectories. To handle the pose drift issue of SLAM and SfM for large-scale scene reconstruction and rendering, Wu et al. [72] introduce bundle adjusting neural radiance field, which jointly optimizes camera pose & scene representation, achieving promising results in both indoor and outdoor scenes.
Hybrid Scene Representations. Based on the observation that a sparse point cloud of a scene is often easily obtained by multi-view stereo, PointNeRF [76], PointNeRF++ [64], and TetraNeRF [28] are proposed to combine the merits between explicit scene geometry and implicit neural representations. They have demonstrated impressive visual quality and high rendering speed compared to previous pure implicit neural representations. To handle asynchronously captured LiDAR data and exposure variation between captured images in large urban outdoor environments, Rematas et al. [53] present Urban Radiance Field (URF), which also leverages segmentation maps to supervise densities on rays pointing at the sky.
Utilizing neural networks for scene rendering often comes with a significant trade-off in rendering speed and quality. To address this, the Plenoxels [14] achieves a remarkable two orders of magnitude acceleration in photorealistic view synthesis speed compared to Neural Radiance Fields, maintaining visual quality through optimization of a sparse 3D grid with spherical harmonics from calibrated images. Extending this concept, Wang et al. [69] combine Fourier and hyperspherical harmonics, achieving dynamic reconstruction. Chen et al. introduce TensoRF [8], a groundbreaking approach to radiance field modeling that represents scenes as 4D tensors, leveraging CP and VM decompositions for improved rendering quality and reduced memory footprint compared to NeRF. Building upon this innovation, Jin et al. [23] propose TensoIR, an inverse rendering method that combines tensor factorization and neural fields for efficient estimation of scene geometry, surface reflectance, and environment illumination.
@@
Back to Explicit. Recently, 3D Gaussian splitting (3DGS) [25] revolutionizes the view synthesis quality for unbounded and complete scenes with 1080p resolution rendering. It represents sparse points from SfM/SLAM/LiDAR sensors with 3D Gaussians and directly optimizes the position, anisotropic covariance, opacity, and spherical harmonic (SH) coefficients of each Gaussian. Benefitting from a tile-based rasterizer solution, 3DGS achieves super-fast and ultra-fine rendering quality. Our method is built upon the 3DGS technique and combines the merits of conventional explicit point cloud/mesh and Gaussian representation for large-scale scene reconstruction and rendering.

### v2

@@ Abstract
Large garages are ubiquitous yet intricate scenes that present unique challenges due to their monotonous colors, repetitive patterns, reflective surfaces, and transparent vehicle glass. Conventional Structure from Motion (SfM) methods for camera pose estimation and 3D reconstruction often fail in these environments due to poor correspondence construction. To address these challenges, we introduce LetsGo, a LiDAR-assisted Gaussian splatting framework for large-scale garage modeling and rendering. We develop a handheld scanner, Polar, equipped with IMU, LiDAR, and a fisheye camera, to facilitate accurate data acquisition. Using this Polar device, we present the GarageWorld dataset, consisting of eight expansive garage scenes with diverse geometric structures, which will be made publicly available for further research. Our approach demonstrates that LiDAR point clouds collected by the Polar device significantly enhance a suite of 3D Gaussian splatting algorithms for garage scene modeling and rendering. We introduce a novel depth regularizer that effectively eliminates floating artifacts in rendered images. Additionally, we propose a multi-resolution 3D Gaussian representation designed for Level-of-Detail (LOD) rendering. This includes adapted scaling factors for individual levels and a random-resolution-level training scheme to optimize the Gaussians across different resolutions. This representation enables efficient rendering of large-scale garage scenes on lightweight devices via a web-based renderer. Experimental results on our GarageWorld dataset, as well as on ScanNet++ and KITTI-360, demonstrate the superiority of our method in terms of rendering quality and resource efficiency.
Figure 1: We present LetsGo - an explicit and efficient end-to-end framework for high-fidelity rendering of large-scale garages. We design a handheld Polar scanner to capture RGBD data of expansive parking environments and have scanned a garage dataset, named GarageWorld, comprising eight garages with different structures. Our LiDAR-assisted Gaussian primitives approach along with GarageWorld dataset enables various applications, such as autonomous vehicle localization, navigation and parking, as well as VFX production.
**footnotetext: Authors contributed equally to this work.footnotetext: These authors are corresponding authors.
@@ 1 Introduction
Modeling garage environments accurately is crucial for various applications such as autonomous vehicle testing, architectural planning, and game design. Garages present a unique set of challenges due to their complex geometries, varying lighting conditions, and frequent presence of obstacles. The ability to create detailed and realistic 3D models of garages can significantly enhance the effectiveness of these applications, providing more accurate simulations and analyses. For visual artists, garages represent a frontier in visual simulation that merges the aesthetic with the technical, offering a canvas where the intricacies of light, shadow, texture, and space coalesce. The complex interplay of artificial and natural lighting within the confines of a garage, with its reflective surfaces, varying materials, and intricate geometries, provides a rigorous testbed for modeling and rendering, pushing the boundaries of what is achievable in virtual environments.
@@
However, capturing and rendering garage environments pose significant difficulties. These spaces often have low-light conditions, textureless surfaces, and a high degree of clutter (Fig.5). The intricacies of garage layouts, including narrow spaces and reflective surfaces, further complicate the modeling process. Garages with internal circular or spiral paths often lead to incomplete data and ambiguities in spatial relationships. Additionally, their extensive spatial area underscores the necessity for efficient 3D representations and lightweight rendering techniques to facilitate real-time interaction and visualization, particularly for scenarios demanding rapid situational assessment, such as navigation and path adjustment.
@@
Existing methods for garage modeling typically rely on either manual measurements or conventional photogrammetry and LiDAR scanning. Traditional computer vision techniques, such as Structure from Motion (SfM)[71, 65] and Multi-view Stereo (MVS[91, 24]), often struggle in these environments due to the prevalence of texture-less regions and repetitive structural designs. These methods frequently fail to extract sufficient feature points and establish accurate feature correspondences necessary for estimating camera poses. Active sensing technologies based on LiDAR can calculate camera poses and scene geometry using SLAM algorithms, but the reflective materials and transparent car windows common in garages lead to geometric inaccuracies. Moreover, LiDAR data tend to be sparse, containing many holes that corrupt high-frequency textures essential for rendering the color appearance of the scene.
@@
While recent advances in neural representation, particularly Neural Radiance Fields (NeRF)[51], have shown promise in producing high-quality renderings, they come with high computational costs and lengthy training times. Although enhancements [72, 95, 88, 16] to NeRF aim to optimize training duration and visual rendering quality, integrating such implicit representations into conventional graphics rendering pipelines and tools for rapid 3D content applications remains challenging. The emerging 3D Gaussian Splatting (3DGS)[34] method revisits explicit representations, using 3D Gaussians to articulate the geometry and appearance of scenes, achieving high-quality scene modeling and rendering. Recent works [35, 47, 61, 48, 70] extend the 3DGS approach to model large-scale outdoor scenes, achieving impressive results. However, few methods address the unique challenges of modeling large-scale indoor scenes like underground garages, where low lighting, large texture-less regions, and repetitive patterns complicate the establishment of sufficient feature correspondences between different images.
@@
This paper introduces LetsGo, an explicit and efficient end-to-end modeling scheme for high-fidelity rendering of large-scale garages. Our key innovation is the integration of calibrated LiDAR points into 3D Gaussian splatting algorithms. We design a handheld Polar scanner, which combines IMU, LiDAR, and a fisheye camera for robust relative pose estimation, specifically tailored for expansive garage data collection. We scan eight large-scale garages, collectively named GarageWorld, using this Polar scanner. To our knowledge, this dataset is the first of its kind aimed at large-scale garages and will be made available to the community. Our experiments demonstrate that LiDAR points collected by the Polar device effectively support various Gaussian splatting algorithms for detailed garage scene representation. To enhance the quality of 3D Gaussian rendering, we introduce a depth regularizer that uses depth priors as supervisory signals, significantly reducing floating artifacts and enabling high rendering quality.
As the scene size increases, the memory demands for rendering large amounts of 3D Gaussians can exceed the capabilities of even high-end GPUs. To address this challenge, we propose a multi-resolution 3D Gaussian representation tailored for Level-of-Detail (LOD) rendering. This approach dynamically adjusts based on the camera’s position, orientation, and viewing frustum, allowing for real-time, high-quality rendering of expansive scenes. Specifically, we construct different levels of Gaussians at varying resolutions, where lower-resolution levels capture coarse scene characteristics and higher-resolution levels reconstruct fine, high-frequency details. We employ tailored scaling factors for each level and a random-resolution-level training scheme to optimize the Gaussians across different levels.
During rendering, we introduce a novel level selection strategy that optimizes the trade-off between visual fidelity and device performance by considering the distance between 3D Gaussians and the rendering viewpoint. With our multi-resolution 3D Gaussian framework, we develop an LOD PC viewer that achieves rendering speeds up to four times faster than traditional 3DGS viewers on high-performance GPUs (e.g., RTX 3090). Additionally, we offer a lightweight web renderer designed to support LOD rendering across various consumer-level devices, including laptops and tablets. We will release our source codes, including our training code, high-performance PC viewer, and lightweight web viewer, to facilitate reproducible research.
@@
Our results, gathered from the GarageWorld dataset as well as ScanNet++[13] and KITTI-360[44] datasets, indicate that our approach not only surpasses other methods in rendering quality but also maintains high rendering efficiency. The GarageWorld dataset and the LetsGo framework for large-scale garage modeling and rendering enable various applications, including autonomous driving, localization, navigation, visual effects, etc.
@@ 2 Related Work
Conventional algorithms for reconstructing large-scale scenes include Structure from Motion (SfM) [82, 74, 52], Simultaneous Localization and Mapping (SLAM) [40, 4, 5, 6] and Multi-View Stereo (MVS) [68, 25, 23]. They are dedicated to discerning the three-dimensional structure of a scene through the sequential or multi-view analysis of two-dimensional image frames. All these methods leverage feature tracking and multi-view consistency to recover the 3D scene structures. SfM- and SLAM-based methods [12, 1, 19, 28, 78, 64, 42] estimate poses of input images and recover the scene structure jointly. However, their main purpose is pose estimation, and the recovered scene point clouds are always sparse, making it difficult for high-quality free-view synthesis. While MVS-based methods [41, 15, 29, 90, 87, 86], especially deep-based approaches, compute a dense depth map for each input image, the constructed scenes often lack accuracy and robustness in texture-less and complex scenes.
NeRF and 3D Gaussian Splatting Variants.
Recent advances [55, 9, 93, 36, 58, 98, 97] in neural scene representation have significantly impacted novel view synthesis. Neural Radiance Fields (NeRF), introduced by Mildenhall et al. [51], have revolutionized 3D reconstruction with a novel framework for detailed scene capture. Subsequent improvements [73, 38, 62, 60] have aimed at enhancing the visual fidelity and computational efficiency of NeRF. MipNeRF [2] and MipNeRF360 [3] address aliasing via a novel conical frustum rendering technique and a nonlinear scene representation, respectively. Techniques such as Plexnoxels [20], Instant-NGP [53], and TensoRF [8] have expedited rendering by integrating explicit encoding methods with compact MLP networks. Furthermore, Block-NeRF [75] and Mega-NeRF [79] facilitate the application of NeRF to extensive scenes. F2-NeRF [81] introduces a space-warping method for handling arbitrary camera trajectories. ScaNeRF [84] optimizes camera pose and scene representation jointly to address pose drift in large-scale scene reconstruction. Unlike these NeRF-based approaches, our method employs an explicit 3D Gaussian representation with depth priors from our RGBD scanner, enhancing realism and efficiency.
3D Gaussian Splatting (3DGS)[34] is a transformative approach characterized by its efficiency and lifelike visual quality. Extensive research [48, 54, 33, 49, 7, 83, 31, 77] has built on 3DGS. Mip-Splatting [94] mitigates artifacts from varying view sampling rates, and GaussianPro [10], which enhances Gaussian distribution with 2D image constraints. DN-Splatter [80] and DNGaussian [43] use depth and normal cues for refinement. SuGaR [27] introduces Gaussian alignment regularization for explicit mesh extraction, and GaussianSurfels [14] flattens 3D Gaussian ellipsoids into 2D ellipses for accurate surface reconstructions. 2DGS [30] transforms 3D volumes into 2D planar Gaussian disks, employing a perspective-accurate 2D splatting process aligned with geometric surfaces. Despite these advances focusing on small-scale or object-level scenes, our approach innovates with a depth regularizer and a multi-resolution Gaussian representation designed for large-scale scenes, achieving realistic visual quality and enhanced rendering efficiency.
Gaussian Splatting for Large-Scale Scenes.
To effectively manage large-scale scenes, VastGaussian [46] explores various partitioning strategies and introduces an appearance embedding module to enhance Gaussian training. Concurrent works such as DrivingGaussian [99] and Street Gaussian [89] adapt Gaussian methodologies to the dynamic contexts of urban landscapes and autonomous driving scenarios, respectively. In addition, multiple studies are integrating the Level of Detail (LOD) rendering technique with 3D Gaussian Splatting to balance rendering quality and speed. Traditional mesh-based LOD strategies [57, 11] create varying levels of detail in object meshes, selecting the appropriate level based on the viewer’s distance to the target or the desired quality. Point-based LOD methods, like Potree [66] and FastLOD [67], are designed to efficiently render large 3D point clouds, which is particularly beneficial for LiDAR data visualization.
Recently, Octree-GS [61] proposes representing Gaussians within an octree structure to render fine details at different viewing scales. CityGaussian [47] employs an LOD strategy for efficiently training and rendering of large-scale 3DGS. Hierarchical 3DGS [35] introduces a hierarchical 3DGS, showing promising results in rendering extensive walk-through datasets. However, these methods lack of on-demand rendering schemes and require high-computational devices to load complete scenes for real-time rendering, making them unsuitable for lightweight devices such as tablets or laptops. In contrast, our LetsGo method employs a multi-resolution Gaussian representation coupled with an on-demand rendering scheme. This approach achieves rendering speeds four times faster than the original 3DGS and ensures compatibility with lightweight devices, facilitating real-time rendering in large-scale garage scenes.
@@
Table 1: Detailed illustrations of our GarageWorld dataset, including various datasets categorized by their environment type, geometric features, area size, number of images, point count, face count, and lighting conditions
Campus 1 Underground Flat & Sloped Paths 38447.86 8479 1.9B 121.9M Uniform Lighting
Campus 2 Underground Flat & Sloped Paths 28046.37 7772 1.4B 95.8M Uniform Lighting
Shopping Mall 1 Indoor (Multi-floors) Spiral & Circular Paths 32646.68 5792 1.14B 40.8M Uneven Lighting
Shopping Mall 2 Outdoor Spiral & Sloped Paths 13495.92 2280 0.6B 77.1M Natural Lighting
Shopping Mall 3 Underground Flat & Sloped Paths 30246.07 13296 1.6B 106.1M Uniform Lighting
Office Building Underground With Mechanical Parking System 22159.25 9308 1.15B 72.3M Motion Sensor Lighting
Arts Center Underground Flat & Sloped Paths 10392.32 5779 0.52B 31.4M Uniform Lighting
Subway Garage Underground Flat & Sloped Paths 7109.12 3607 0.70B 43.5M Uniform Lighting